OpenStudy (anonymous):

If the problem type is labeled “ClickAndPlace”, click the item to select it and then click the location where it belongs; otherwise, click the item and drag it into place. Express answer in exact form. Find the area of the larger segment whose chord is 8" long in a circle with an 8" radius. help me

OpenStudy (anonymous):

come to this web page. I will explain you there. http://www.twiddla.com/583902

OpenStudy (anonymous):

Find the total area of the circle: pi*r^2 = pi*8^2= 201.062 in^2 The chord and the two radii form an equilateral triangle, all angles 60 degrees. Find the area of the portion contained in the 60 degree area 60* 201.062/360 = 33.51 in^2 Find the area of the equilateral triangle. = 27.71 in^2 Find the area of chords smaller segment: 33.51 - 27.71 = 5.8 in^2 Find the area of the larger segment: 201.06 - 5.8 = 195.26 in^2